# Consciousness-Inspired Spatio-Temporal Abstraction for Better Generalization

### 04 Feb 2024 v3

@@ 6 Conclusion & Future Work
A.1 Skipper-once Scalability
@@
A.2 Skipper-regen Scalability
@@
A.3 modelfree Baseline Scalability
@@ A.5 Director Scalability
A.6 Comparative Generalization Performance on Different Numbers of Training Tasks
The comparative results of all agents performing on each training configurations, i.e. different numbers of training tasks, are presented in Fig. 10, Fig. 11, Fig. 12, Fig. 13, Fig. 14 and Fig. 15.
@@ A.6 Comparative Generalization Performance on Different Numbers of Training Tasks
A.6.1 Statistical Significance & Power Analyses
Besides visually observing generally non-overlapping confidence intervals, we present the pairwise $t$-test results of Skipper-once and Skipper-regen against the other methods, as well as, if significant, power analyses to determine if the number of seed runs ($20$) was enough to make the significance claim. These results are shown in Tab. 1 and Tab. 2, respectively.
@@ A.6.1 Statistical Significance & Power Analyses
$t$-test significance is set to be $0.05$.
Effect size equals to the difference of the means of the compared pairs (Colas et al., 2018).
For significant comparisons, the minimum number of seeds to achieve statistical power $0.2$ is provided.
Cells are marked bold if results not significant or not enough seeds to achieve statistical power.
@@
$t$-test significance is set to be $0.05$.
Effect size equals to the difference of the means of the compared pairs (Colas et al., 2018).
For significant comparisons, the minimum number of seeds to achieve statistical power $0.2$ is provided.
Cells are marked bold if results not significant or not enough seeds to achieve statistical power.
Appendix B Ablation & Sensitivity
B.1 Validation of Effectiveness on Stochastic Environments
We present the performance of the agents in stochastic variants of the used environment. Specifically, in these tasks, with probability $0.1$ where each action an agent takes could be changed into a random action. We present the $50$-training tasks performance evolution in Fig. 16. The results validate the compatibility of our agents with stochasticity in environmental dynamics. Notably, the performance of the baseline deteriorated to worse than even Director with the injected stochasticity. The compatibility of Hierarchical RL frameworks to stochasticity has been investigated in Hogg et al. (2009).
@@ B.1 Validation of Effectiveness on Stochastic Environments
B.2 Ablation for Spatial Abstraction
We present in Fig. 17 the ablation results on the spatial abstraction component with Skipper-once agent, trained with $50$ tasks. The alternative component of the attention-based bottleneck, which is without the spatial abstraction, is an MLP on a flattened full state. The results confirm significant advantage in terms of generalization performance as well as learning speed, brought by the introduced spatial abstraction technique.
@@ B.2 Ablation for Spatial Abstraction
B.3 Accuracy of Proxy Problems & Checkpoint Policies
We present in Fig. 18 the ablation test results on the accuracy of proxy problems as well as the checkpoint policies of the Skipper-once agents, trained with $50$ tasks. The ground truths are computed via Dynamic Programming (DP) on the optimal policies, which are also suggested by DP. Concurring with our theoretical analyses, the results indicate that the performance of Skipper is determined (bottlenecked) by the accuracy of the proxy problem estimation on the high-level and the optimality of the checkpoint policy on the lower level. Specifically, the curves for the generalization performance across training tasks, as in (a) of 18, indicate that the lower than expected performance is a composite outcome of errors in the two levels. In the next part, we address a major misbehavior of inaccurate proxy problem estimation - chasing delusional targets.
@@ B.3 Accuracy of Proxy Problems & Checkpoint Policies
Figure 18: Skipper-once Empirical Performance v.s. ground truths: both the optimal policy and optimal plan variants are calculated via DP on the environment dynamics. The default deterministic setting induces the fact that combining optimal policy and optimal plan results in $1.0$ success rate. The figures suggest that the learned agent is limited by errors both in the proxy problem estimation and the checkpoint policy $\pi$. Each agent is trained with $50$ environments and each curve is processed from $20$ independent seed runs.
B.4 Training Initialization: uniform v.s. same as evaluation
We present in Fig. 19 the comparative results on the training setting: whether to use uniform initial state distribution or not. The non-uniform starting state distributions introduce additional difficulties in terms of exploration and therefore globally slow down the learning process. These results are obtained from training on $50$ tasks. We conclude that given similar computational budget, using non-uniform initialization only slows down the learning curves, and thus we use the ones with uniform initialization for presentation in the main paper.
@@
B.5 Ablation: Vertex Pruning
As mentioned previously, each proxy problem in the experiments are reduced from $32$ vertices to $12$ with such techniques. We present the comparative performance curves of the used configuration against a baseline that generates $12$-vertex proxy problems without pruning. We present in Fig. 20 these ablation results on the component of $k$-medoids checkpoint pruning. We observe that the pruning not only increases the generalization but also the stability of performance.
@@ B.5 Ablation: Vertex Pruning
B.6 Sensitivity: Number of Vertices
We provide a sensitivity analysis to the number of checkpoints (number of vertices) in each proxy problem. We present the results of Skipper-once on $50$ training tasks with different numbers of post-pruning checkpoints (all reduced from $32$ by pruning), in Fig. 21. From the results, we can see that as long as the number of checkpoints is above $6$, Skipper exhibits good performance. We therefore chose $12$, the one with minimal computation cost, as the default hyperparameter.
@@ B.6 Sensitivity: Number of Vertices
B.7 Ablation: Planning over Proxy Problems
We provide additional results to intuitively understand the effectiveness of planning over proxy problems. This is done by comparing the results of Skipper-once with a baseline Skipper-goal that blindly selects the task goal as its target all the time. We present the results based on $50$ training tasks in Fig. 22. Concurring with our vision on temporal abstraction, we can see that solving more manageable sub-problems leads to faster convergence. The Skipper-goal variant catches up later when the policy slowly improves to be capable of solving long distance navigation.
@@ B.7 Ablation: Planning over Proxy Problems
train the distance estimator $D$ as if these are sampled from the HER
Algorithm 1 Delusion Suppression
Appendix C Delusion Suppression
RL agents are prone to blindly optimizing for an intrinsic objective without fully understanding the consequences of its actions. Particularly in model-based RL or in Hierarchical RL (HRL), there is a significant risk posed by the agents trying to achieve delusional future states that do not exist within the safety constraints. With a use of a learned generative model, as in the proposed framework, such risk is almost inevitable, because of uncontrollable generalization effects.
Generalization abilities of the generative models are a double-edged sword. The agent would take advantage of its potentials to propose novel checkpoints to improve its behavior, but is also at risk of wanting to achieve non-existent unknown consequences. In our framework, checkpoints proposed by the generative model could correspond to non-existent “states” that would lead to delusional edge estimates and therefore confuse planning. For instance, arbitrarily sampling partial descriptions may result in a delusional state where the agent is in a cell that can never be reached from the initial states. Since such states do not exist in the experience replay, the estimators will have not learned how to handle them appropriately when encountered in the generated proxy problem during decision time. We present a resulting failure mode in Fig. 23.
(a) generalization performance across training tasks
(b) distance estimation error (L1) for delusional edges
(c) frequency of selecting delusional targets
(d) target optimality
Figure 24: Ablation Results for Skipper-once with the proposed Delusion Suppression Technique: each curve and corresponding error bar (95% CI) are processed from $20$ independent seed runs. a) the performance across training tasks is shown. A more optimal performance can be achieved with Skipper-once in training tasks, when delusions are suppressed; b) During training interactions, the error in estimated (truncated) distance from and to delusional targets are significantly reduced with the technique; c) The frequency of selecting a delusional target is reduced to almost negligible during the whole training process; d) The optimality of target checkpoint during training can be improved by the suppression. Each agent is trained with $50$ environments and each curve is processed from $20$ independent seed runs.
The delusion suppression technique can also be used to help us understand the failure modes of LEAP, which can be found in the later contents.
Appendix D Recovering Discounts
We can recover the distribution of the cumulative discount by replacing the support of the discretized truncated distances with the corresponding discounts, as shown in Fig. 25. Specifically, the problem we wanted to dodge was $\mathbb{E}[\gamma^{D}]\neq\gamma^{\mathbb{E}[D]}$. Luckily, the probability of having a trajectory length of 4 under policy $\pi$ from state $s_{t}$ to $s_{\odot}$ is the same as a trajectory having discount $\gamma^{4}$. The estimated distribution over distances is used to recover on a different support the corresponding distribution of discounts.
Appendix E Skipper Implementation Details
The PyTorch-based source code of experiments is uploaded in the supplementary materials, where reviewers could find the detailed architectures that may be difficult to understand from the following descriptions. The hyperparameters introduced by Skipper can be located in Alg. 2.
The agent is based on a distributional prioritized double DQN. All the trainable parameters are optimized with Adam at a rate of $2.5\times 10^{-4}$ (Kingma & Ba, 2014), with a gradient clipping by value (maximum absolute value $1.0$). The priorities for experience replay sampling are equal to the per-sample training loss, which is a sum of all KL-divergences of the estimators (cross-entropy for the termination estimator), and the ELBO loss of the checkpoint generator.
E.1 Full State Encoder
E.2 Partial State Selector
The selector $\sigma$ is implemented with one-head (not multiheaded, therefore the output linear transformation of the default multihead attention implementation in PyTorch is disabled.) top-$4$ attention, with each local perceptive field of size $8\times 8$ cells. Layer normalization (Ba et al., 2016) is inserted before and after the spatial abstraction.
E.3 Estimators
The estimators, which operate on the partial states, are $3$-layered MLPs with $256$ hidden units. An additional estimator for termination is learned, which instead of taking a pair of partial state input, takes only one, and is learned to classify terminal states with cross-entropy loss. The distance from terminal states to other states would be overwritten with $\infty$. The internal $\gamma$ for intrinsic reward of $\pi$ is $0.95$, while the task $\gamma$ is $0.99$
The estimators use distributional TD learning (Dabney et al., 2018). Instead of letting the value estimator output a scalar, we ask it to output a histogram (taking a softmax over a vector output). We regress the histogram towards the TD target, where such target is converted into a skewed histogram, towards which KL-divergence can be used to accomplish the estimation. At the output, there are $16$ bins for each histogram estimation (value, reward, distance).
E.4 Checkpoint Generator
The checkpoint generator is implemented as follows:
  • The generator operates on observation inputs and outputs, because of its compactness and the equivalence to full states under full observability in our experiments;
  • The context extractor $\mathcal{E}_{c}$ is a $32$-dimensional BabyAI embedder. It encodes an input observation into a representation of the episodic context;
  • The partial description extractor $\mathcal{E}_{z}$ is made of a $32$-dimensional BabyAI embedder, followed by $3$ aforementioned residual blocks with $3\times 3$ convolutions (doubling the feature dimension every time) in between, followed by global maxpool and a final linear projection to the latent weights. The partial descriptions are $6$ binary latents that each could represent $64$ different checkpoints. Similar to VQ-VAE (Van Den Oord et al., 2017), we use the argmax of the latent weights as partial descriptions, instead of sampling according to the softmax-ed weights. This enables easy comparison of current state to the checkpoints in the partial description space, because each state deterministically corresponds to one partial description. In our implementation, we identify reaching a target checkpoint if the partial description of the current state matches that of the target.
  • The whole checkpoint generator is trained end-to-end with a standard VAE loss. That is the sum of a KL-divergence for the agent’s location, and the entropy of partial descriptions, weighted by $2.5\times 10^{-4}$, as suggested in https://github.com/AntixK/PyTorch-VAE. Note that the per-sample losses in the batches are not weighted according to priority from the experience replay.
As in our experiments, if one does not want to generate non-goal terminal states as checkpoints, we could also seek to train on reversed $\langle S^{\odot},S_{t}\rangle$ pairs. In this case, the checkpoints to reconstruct will never be terminal.
E.5 HER
Each agent-environment transition is further duplicated into $4$ hindsight transitions at the end of each episode. Each of these transitions is combined with a randomly sampled observation from the same trajectory as the relabelled ”goal”. The size of the hindsight buffer is extended to $4$ times that of the baseline that does not learn from hindsight accordingly, that is, $4\times 10^{6}$.
E.6 Planning
As introduced, we use value iteration over options (Sutton et al., 1999) to plan over the proxy problem represented as an SMDP. We use the matrix form $Q=R_{S\times S}+\Gamma V$, where $R$ and $\Gamma$ are the estimated edge matrices for cumulative rewards, respectively. Note that this notation is different from the ones we used in the manuscript. The checkpoint value $V$, initialized as all-zero, is taken on the maximum of $Q$ along the checkpoint target (the actions for $\mu$) dimension. At decision time, we run each planning for $5$ iterations. The edges from the current state towards other states are always set to be one-directional, and the self-loops are also deleted from the graph. This means the first column as well as the diagonal elements of $R$ and $\Gamma$ are all zeros. Besides pruning edges based on the distance threshold, as introduced in the main paper, the terminal estimator is also used to prune the matrices: the rows corresponding to the terminal states are all zeros.
E.7 Hyperparameter Tuning
This leads to how we tuned the pruning (distance) threshold: we fully used the advantage of our experiments on a fully DP-solvable environment: with a snapshot of the agent during its training, we can sample many ¡starting state, target state¿ pairs and calculate the ground truth distance between the pair, as well as the failure rate of reaching from the starting state to the target state given the current policy $\pi$, which is again enabled by DP, then plot them as the $x$ and $y$ values respectively for visualization. We found such curves to evolve from high failure rate at the beginning, to a monotonically increasing curve, where at small true distances, the failure rates are near zero. We picked $8$ because the curve starts to grow explosively when the true distances are more than $9$.
$k$ for $k$-medoids Pruned via the sensitivity analysis provided in the Appendix.
We tuned this by running a sensitivity analysis on Skipper agents with different $k$’s, whose results are presented previously in this Appendix,.
Appendix F LEAP Implementation Details
The LEAP baseline has been implemented from scratch for our experiments, since the original open-sourced implementation [^4] https://github.com/snasiriany/leap was not compatible with environments with discrete action spaces. LEAP’s training involves two pretraining stages, that are, generator pretraining and distance estimator pretraining, which are named the VAE and RL pretraining originally. Despite our best effort, that is to be covered in details, we found that LEAP was unable to get a reasonable performance in its original form after rebasing on a discrete model-free RL baseline.
We tried to identify the reasons why the generalization performance of the adapted LEAP was unsatisfactory: we found that the original VAE used in LEAP is not capable to handle even few training tasks, let alone generalize well to the evaluation tasks. Even by combining the idea of the context / partial description split (still with continuous latents), during decision time, the planning results given by the evolutionary algorithm (Cross Entropy Method, CEM, Rubinstein (1997a)) almost always produce delusional plans that are catastrophic in terms of performance. This was why we switched into LEAP the same conditional generator we proposed in the paper, and adapted the CEM accordingly, due to the change from continuous latents to discrete.
The original distance estimator based on Temporal Difference Models (TDM) also does not show capable performance in estimating the length of trajectories, even with the help of a ground truth distance function (calculated with DP). Therefore, we switched to learning the distance estimates with our proposed method. Our distance estimator is not sensitive to the sub-goal time budget as TDM and is hence more versatile in environments like that was used in the main paper, where the trajectory length of each checkpoint transition could highly vary. Like for Skipper, an additional terminal estimator has been learned to make LEAP planning compatible with the terminal lava states. Note that this LEAP variant was trained on the same sampling scheme with the HER, and was marked as ”LEAP” in the main paper.
We also did not find that using the pretrained VAE representation as the state representation during the second stage helped the agent’s performance, as the paper claimed. In fact, the adapted LEAP variant could only achieve decent performance after learning a state representation from scratch in the RL pretraining phase.
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
Appendix G Director Implementation Details
Figure 27: An example for simplified observations for Director.
Adaptation. For our experiments with Director (Hafner et al., 2022), we have used the publicly available code [^5] See https://github.com/danijar/director provided by the authors. Except for a few changes in the parameters, which are depicted in Table 3, we have used the default configuration provided for Atari environments. Note that as the Director version in which the worker receives no task rewards performed very badly in our environment, we have used the version in which the worker receives scaled task rewards (referred to as “Director (worker task reward)” in Hafner et al. (2022)). This agent has also been shown to perform better across various domains in Hafner et al. (2022).
Encoder. Unlike Skipper and LEAP agents, the Director agent receives as input a simplified RGB image of the current state of the environment (see Fig. 27). This is because we found that Director performed better with its original architecture, which was designed for image-based observations. To simplify the representation learning of Director as much as possible, we also used a simplified RGB image, instead of a detailed one (where the lava cells have waves on top etc.).
Parameter
Value
replay_size
2M
replay_chunk
12
imag_horizon
8
env_skill_duration
4
train_skill_duration
4
worker_rews
{extr: 0.5, expl: 0.0, goal: 1.0}
sticky
False
gray
False
Training Performance. We investigated why Director is unable to achieve near optimal training performance in the used environment (Fig. 3). As Director was trained solely on environments where it is able to collect long trajectories to train a good enough recurrent world model (Hafner et al., 2022), we hypothesized that Director may perform better in domains where it is able to interact with the environment through longs trajectories to train a good enough recurrent world model (i.e., the agent does not immediately die as a result of interacting with specific objects in the environment). To test this, we experimented with variants of the used environments, where the lava cells are replaced with wall cells, so the agent does not die upon trying to move towards them (we refer to this environment as the “walled” environment). The corresponding results on $50$ training tasks are depicted in Fig. 28. As can be seen, the Director agent indeed performs better within the training tasks than in the environments with lava.
Generalization Performance. We also investigated why Director is unable to achieve good generalization performance in the used environment (Fig. 3). As Director trains its policies solely from the imagined trajectories predicted by its learned world model, we believe that the low generalization performance is due to Director being unable to learn a good enough world model that generalizes to the evaluation tasks. The generalization performances in both the “walled” and regular environments, depicted in Fig. 28, indeed support this argument. Similar to the main paper, we also present experimental results for how the generalization performance changes with the number of training environments that are used. Results in Fig. 29 show that the number of training environments has no effect on the poor generalization performance of Director.
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
Appendix H Pseudo Code (with Hyper-Parameters)
H.1 Overall Skipper Framework
The pseudocode of Skipper is provided in Alg. 2.
H.2 k-medoids based pruning
for each episode do
       //construct the proxy problem, can be done multiple times in one episode too (as in Skipper-regen)
       generate more than necessary (32) checkpoints by sampling from the partial description latent space given the extracted context from the initial state
       k-medoid pruning based on the distance estimations among all checkpoints (down to $12$ vertices)
       use estimators to construct the pruned graph (including a terminal state estimator to correct the estimates)
       prune edges that are too far-fetched according to distance estimations (threshold set to be $8$, same as replan interval)
       for each agent-environment interaction step do
             if decided to explore (DQN-style annealing $\epsilon$-greedy) then
                   take a random action
            else
                         run value iteration (for $5$ iterations) on the proxy problem, select the target checkpoint
                  follow the action suggested by the checkpoint-achieving policy
            if time to train (every $4$ actions) then
                   sample hindsight transitions and train checkpoint-achieving policy, estimators (including a teriminal state estimator) and checkpoint generator
                   [optional delusion control]: generate imaginary states and use them to train the estimators
            save interaction into the trajectory experience replay
      convert trajectory into hindsight samples in HER (relabel $4$ random states as additional goals)
We present the pseudocode of the modified $k$-medoids algorithm for pruning overcrowded checkpoints in Alg. 3. Note that the actual implementation is parallelized, the simplification for presentation is intended for reader’s understanding. The changes upon the original $k$-medoids algorithm is marked in purple. When $k$-medoids is called after the unpruned graph is constructed, $\mathcal{S}_{\vee}$ is set to be the set containing the goal state only. This is intended to span more uniformly in the state space with checkpoints, taking consideration into the goal. Let the estimated distance matrix be $D$, where each element $d_{i}j$ represents the estimated trajectory length it takes for $\pi$ to fulfill the transition from checkpoint $i$ to checkpoint $j$. Since $k$-medoids cannot handle infinite distances (e.g. from a terminal state to another state), the distance matrix $D$ is truncated, and then we take the elementwise minimum between the truncated $D$ and $D^{T}$ to preserve the one-way distances. The matrix containing the elementwise minimums would be the input of the pruning algorithm.
make sure $\mathcal{S}_{\vee}\subset\mathcal{S}_{\odot}$
repeat
       foreach medoid $M_{j}$ do
      Find the medoid $M_{j}$ with the lowest cost $J_{j}$;
       if $M_{j}$ changes then
             make sure $\mathcal{S}_{\vee}\subset\mathcal{S}_{\odot}$
             Replace $M_{j}$ with the data point in $C_{j}$ that minimizes the total cost;
until Convergence (no cost improvement);
Algorithm 3 Checkpoint Pruning with $k$-medoids
Appendix I Weaknesses and Limitations
We would like to discuss the limitations to the current form of the framework, as well as the design choices that we seek to improve in the future:
  • We generate future checkpoints at random by sampling the partial description space. Despite the post-processing such as pruning, the generated checkpoints do not prioritize on the predictable, important states that matter the most to form a meaningful long-term plan. We would like to continue investigating the possibilities along this line;
  • The current implementation is intended for pixel input fully-observable tasks with discrete state and action spaces. Such a minimalistic form is because we wish to isolate the unwanted challenges from other factors that are not closely related to the idea of this work, as well as to make the agent as generalist as possible. On the matter of the compatibility with continuous spaces, Skipper is naturally compatible with continuous actions spaces and the only thing we will need to do is to replace the baseline agent with a compatible one such as TD3 (Fujimoto et al., 2018); on the other hand, for continuous state spaces, the identification of the achievement of a checkpoint becomes tricky. This is due to the fact that a strict identity between the current state and the target checkpoint may be ever established, we either must adopt a distance measure for approximate state equivalence, or rely on the equivalence of the partial descriptions (which is adopted in the current implementation). We intentionally designed the partial descriptions to be in the form of bundles of binary variables, so that this comparison could be done fast and trivially for any forms of the state space; for partial observability, despite that no recurrent mechanism has been incorporated in the current implementation, the framework is not incompatible. To implement that, we will need to augment the state encoder with recurrent or memory mechanisms and we need to make the checkpoint generator directly work over the learned state representations. We acknowledge that future work is needed to verify Skipper’s performance on the popular partially-observable benchmark suites, which requires the incorporation of components to handle partial observability as well as scaling up the architectures for more expressive power;
  • We are far from understanding the precise boundaries of the motivating theory on proxy problems, since it only indicates performance guarantees on the condition of estimation accuracy, which in turn does not correspond trivially to a set of well-defined problems. We are eager to explore, outside the scope of sparse-reward navigation, how this approach can be used to facilitate better generalization, and at the same time, try to find more powerful theories that guide us better;
Appendix J Theoretical Analyses
J.1 Update Rules for Edge Estimation
First, we show that the update rules proposed in the main paper indeed estimate the desired cumulative discount and reward.
For the cumulative discounted reward random variable:
$=\gamma^{T_{\perp}-t}\mathbb{I}{\{S_{T_{\perp}}=s^{\odot}\}},$ (8)
J.2 Performance Bound
Similarly, we can derive a lower bound:
J.3 No Assumption on Optimality

### 16 Mar 2024 v4

@@ E.3.2 Failure Modes: Bad Generalization, Sensitive to Short Trajectories
F.1 Skipper-once Scalability
@@
F.2 Skipper-regen Scalability
@@
F.3 modelfree Baseline Scalability
@@ F.5 Director Scalability
F.6 Generalization Performance on Different Numbers of Training Tasks
The performance of all agents on all training configurations, i.e. different numbers of training tasks, are presented in Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21 and Fig. 22.
@@ F.6 Generalization Performance on Different Numbers of Training Tasks
F.6.1 Statistical Significance & Power Analyses
Besides visually observing generally non-overlapping confidence intervals, we present the pairwise $t$-test results of Skipper-once and Skipper-regen against the compared methods. In addition, if the advantage is significant, we perform power analyses to determine if the number of seed runs ($20$) was enough to make the significance claim. These results are shown in Tab. 2 and Tab. 3, respectively.
@@ F.6.1 Statistical Significance & Power Analyses
$t$ threshold: $0.05$.
Effect size set to be the difference of the means of the compared pairs (Colas et al., 2018).
Cells are bold if results NOT significant or insufficient seeds for statistical power.
For significant cases, the minimum number of seeds for statistical power $0.2$ is provided.
@@
$t$ threshold: $0.05$.
Effect size set to be the difference of the means of the compared pairs (Colas et al., 2018).
Cells are bold if results NOT significant or insufficient seeds for statistical power.
For significant cases, the minimum number of seeds for statistical power $0.2$ is provided.
Appendix G Ablation & Sensitivity
G.1 Validation of Effectiveness on Stochastic Environments
We present the performance of the agents in stochastic variants of the used environment. Specifically, with probability $0.1$, each action is changed into a random action. We present the $50$-training tasks performance evolution in Fig. 23. The results validate the compatibility of our agents with stochasticity in environmental dynamics. Notably, the performance of the baseline deteriorated to worse than even Director with the injected stochasticity. The compatibility of Hierarchical RL frameworks to stochasticity has been investigated in Hogg et al. (2009).
@@ G.1 Validation of Effectiveness on Stochastic Environments
G.2 Ablation for Spatial Abstraction
We present in Fig. 24 the ablation results on the spatial abstraction component with Skipper-once agent, trained with $50$ tasks. The alternative component of the attention-based bottleneck, which is without the spatial abstraction, is an MLP on a flattened full state. The results confirm significant advantage in terms of generalization performance as well as sample efficiency in training, introduced by spatial abstraction.
@@ G.2 Ablation for Spatial Abstraction
G.3 Accuracy of Proxy Problems & Checkpoint Policies
We present in Fig. 25 the ablation results on the accuracy of proxy problems as well as the checkpoint policies of the Skipper-once agents, trained with $50$ tasks. The ground truths are computed via DP on the optimal policies, which are also suggested by DP. Concurring with our theoretical analyses, the results indicate that the performance of Skipper is determined (bottlenecked) by the accuracy of the proxy problem estimation on the high-level and the optimality of the checkpoint policy on the lower level. Specifically, the curves for the generalization performance across training tasks, as in (a) of 25, indicate that the lower than expected performance is a composite outcome of errors in the two levels. In the next part, we address a major misbehavior of inaccurate proxy problem estimation - chasing delusional targets.
@@ G.3 Accuracy of Proxy Problems & Checkpoint Policies
Figure 25: Skipper-once Empirical Performance v.s. ground truths: both the optimal policy and optimal plan variants are assisted by DP. The default deterministic setting induces the fact that combining optimal policy and optimal plan results in $1.0$ success rate. The figures suggest that the learned agent is limited by errors both in the proxy problem estimation and the checkpoint policy $\pi$. Each agent is trained with $50$ environments and each curve is processed from $20$ independent seed runs.
G.4 Training Initialization: uniform v.s. same as evaluation
We compare the agents’ performance with and without uniform initial state distribution. The non-uniform starting state distributions introduce additional difficulties in terms of exploration. In Presented in Fig. 26, these results are obtained from training on $50$ tasks. We conclude that given similar computational budget, using non-uniform initialization only slows down the learning curves without introducing significant changes to our conclusions, and thus we use the ones with uniform initialization for presentation in the main paper.
@@
G.5 Ablation: Vertex Pruning
As mentioned previously, each proxy problem in the experiments are reduced from $32$ vertices to $12$ with such techniques. We compare the performance curves of the used configuration against a baseline that generates $12$-vertex proxy problems without pruning. We present in Fig. 27 these ablation results on the component of $k$-medoids checkpoint pruning. We observe that the pruning not only increases the generalization but also the stability of performance.
@@ G.5 Ablation: Vertex Pruning
G.6 Sensitivity: Number of Vertices
We provide a sensitivity analysis to the number of checkpoints (number of vertices) in each proxy problem. We present the results of Skipper-once on $50$ training tasks with different numbers of post-pruning checkpoints (all reduced from $32$ by pruning), in Fig. 28. From the results, we can see that as long as the number of checkpoints is above $6$, Skipper exhibits good performance. We therefore chose $12$, the one with a rather small computation cost, as the default hyperparameter.
@@ G.6 Sensitivity: Number of Vertices
G.7 Ablation: Planning over Proxy Problems
We provide additional results for the readers to intuitively understand the effectiveness of planning over proxy problems. This is done by comparing the results of Skipper-once with a baseline Skipper-goal that blindly selects the task goal as its target all the time. We present the results based on $50$ training tasks in Fig. 29. Concurring with our vision on temporal abstraction, we can see that solving more manageable sub-problems leads to faster convergence. The Skipper-goal variant catches up later when the policy slowly improves to be capable of solving longer distance navigation.